OpenStudy (jkbo):

Please help simplify -(3a+4b)+5a - 4b

OpenStudy (anonymous):

-(3a+4b)+5a-4b =-3a-4b+5a-4b =5a-3a-4b-4b =2a-8b

OpenStudy (jkbo):

thank you

OpenStudy (anonymous):

You're welcome.